OpenStudy (anonymous):

Solve: |2t + 5| = 3 {1, 4} {-1, -4} {-1, 4} {1, -4}

OpenStudy (anonymous):

(-1, -4)

OpenStudy (anonymous):

2t+5=3 2t=-2 t=-1 -2t-5=3 2t=-8 t=-4 (-1,-4)

OpenStudy (anonymous):

thanks guys i got it

OpenStudy (anonymous):

XD